Suppose Phoebe plans to invest $1,000. She can earn an effective annual rate of 5% on Security A, while Security B has an effective annual rate of 12%. After 11 years, the compounded value of Security B should be more than twice the compounded value of Security A. (Ignore risk, and assume that compounding occurs annually.) SOLVE USING FORMULA

A). TRUE

B). FALSE
Answer #1

Security A value at the end of 11 years = 1000 * (1 + R)^N where N = 11 and R = 5%

= 1000 * (1+5%)^11 = 1710.34

Security B value at the end of 11 years = 1000 * (1 + R)^N where N = 11 and R = 12%

= 1000 * (1+12%)^11 = 3478.55

Security A double price = 1710.34 *2 = 3420.68

Thus the value of Security B is more than twice the value of Security A

The statement is TRUE